OpenStudy (anonymous):

Determine if the following is a linear transformation. If it is, find the standard matrix of transformation. T:P1-->P2 such that T(p(t)) = integral of p(t)dt from 0 to x. (I don't know how to make an integral sign on this. sorry!)

OpenStudy (anonymous):

\[\int\limits_{0}^{x} p(t) dt\]

OpenStudy (helder_edwin):

do u know the two basic properties of integrals?

OpenStudy (anonymous):

In order for it to be a linear transformation T(V1+V2)= T(V1) + T(V2) and T(kV1)=k T(V1). Right?

OpenStudy (helder_edwin):

\[ \large \int(A+B)=\int A+\int B \] and \[ \large \int(\alpha K)=\alpha\int K \]

OpenStudy (anonymous):

Oh, those rules. Yes I am familiar with them. I misinterpreted what you were asking for.

OpenStudy (helder_edwin):

so for instance \[ \large T(p_1(t)+p_2(t))=\int_0^x[p_1(t)+p_2(t)]\,dt \] \[ \large =\int_0^xp_1(t)\,dt+\int_0^xp_2(t)\,dt=T(p_1(t))+T(p_2(t)) \]

OpenStudy (anonymous):

Alright, I follow you so far. How would go about finding out if it was a linear transformation?

OpenStudy (helder_edwin):

one step missing. \[ \large T(\alpha p_1(t))= \]

OpenStudy (anonymous):

Ok, and those steps illustrate that it is a linear transformation?

OpenStudy (helder_edwin):

they don't "illustrate" they demonstrate that it IS a linear transformation.

OpenStudy (helder_edwin):

complete the second step. to see if u got it.

OpenStudy (anonymous):

\[\alpha\] is a constant so it can be moved out side leaving you with \[\alpha \]\[P _{1}(t)\] which demonstrates that it is a linear transformation.

OpenStudy (helder_edwin):

\[ \large T(\alpha p_1(t))=\int_0^x[\alpha p_1(t)]\,dt= \alpha\int_0^xp_1(t)\,dt=\alpha\cdot T(p_1(t)) \]

OpenStudy (helder_edwin):

now. u have demonstrated that T is in fact a linear transformation.

OpenStudy (anonymous):

ok, that part of the problem makes it so much more sense now! Thank you! to find the standard matrix of the transformation, would you simply use the basis {1, t, \[t^{2}\] }?

OpenStudy (helder_edwin):

yes. i guess that is what it means.

OpenStudy (anonymous):

ok. thank you so much for your help!

OpenStudy (helder_edwin):

u have the following \[ \large P_1=\langle1,x\rangle \] and \[ \large P_2=\langle1,x,x^2\rangle \]

OpenStudy (helder_edwin):

right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

sorry. a typo (considering the wording of your problem) \[ \large P_1=\langle1,t\rangle \] and \[ \large P_2=\langle1,t,t^2\rangle \]

OpenStudy (helder_edwin):

now compute \[ \large T(1)= \] and \[ \large T(t)= \]

OpenStudy (anonymous):

I feel like I'm making an error (big surprise) I get T(1)=x and T(t)=\[x^{2}\]

OpenStudy (helder_edwin):

check the second one.

OpenStudy (anonymous):

T(t)=\[\frac{ x^{2} }{ 2 }\]?

OpenStudy (helder_edwin):

great.

OpenStudy (helder_edwin):

so now you have: \[ \large T(1)=x=\color{red}{0}\cdot1+\color{red}{1}\cdot x+\color{red}{0}\cdot x^2 \] and \[ \large T(t)=\frac{x^2}{2}=\color{red}{0}\cdot1+\color{red}{0}\cdot x +\color{red}{\frac{1}{2}}\cdot x^2 \] right?

OpenStudy (helder_edwin):

can u write the matrix from this?

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 0 \\ 1 & 0 \\ 0 & 1/2\end{matrix}\right]\]?

OpenStudy (helder_edwin):

yes. good work.

OpenStudy (anonymous):

Just to clarify, that is the standard matrix of the transformation?

OpenStudy (helder_edwin):

yes. because u used the standard basis of each space.

OpenStudy (anonymous):

I understand this problem so much better now. Thank you so much for your help. I really appreciate it!

OpenStudy (helder_edwin):

u r very welcome